### v4

@@ 3.1.2 Annotating Struggle
Since the original videos are trimmed uniformly into 10-second segments we focus on ‘low-level’ struggling cues which we believe to be the most transferable between participants and scenarios. Struggle markers may include shaky hands, pausing actions, dropping items, or action repetition. While there is a variety of ways one might ask an observer to estimate struggle, we chose to ask observers to rate video segments using a four-point scale varying from definitely non-struggle to definitely struggle. Table 1 lists the descriptions for each struggle level. The value of a finer, four-point gradation of struggle supports the longer-term goal of assistive systems. For example, the support provided to a person who is ‘slightly’ struggling may consist of a few basic instructions or reminders. However, a person determined to be definitely struggling may require more detailed guidance or even recommendations to stop the task to prevent injury.
This scoring approach allows naive observers to indicate the degree of struggle/non-struggle. Note, we forced annotators to choose between struggle and non-struggle, instead of allowing a neutral option. This avoided centring strategies on ambiguous judgements and allowed easy binarisation of annotations during early pilot approaches that classed only struggle presence/absence. Throughout our research, we found that indicating the degree of struggle is necessary to address significant challenges in classifying fine-grained features.
As self-reporting levels of struggle can be unreliable and provide only a single judgement, we focus on a balance between expert and naive observers. We collected two distinct types of struggle determination annotations: naive ‘voter labels’ obtained from Amazon Mechanical Turk (MTurk) participants, and ‘expert annotations’ (or ‘golden annotation’ (GA)) obtained from the author who collected the video dataset, who repeatedly viewed them during analysis, and annotated action and error labels in the EPIC-Tent dataset. For each 10s video clip, there is only one expert annotation label, and there are multiple voters’ labels that form a frequency distribution. This distribution consists of 15 or 20 voter labels. To extract a single struggle label from this distribution, we take the mode. Since voter labels and expert annotation come from distinct groups of individuals with differing prior knowledge, label distributions can vary across all video segments for each dataset. Consistency between the voters’ modes and expert annotations indicates that the struggle-level annotations of the videos are more reliable. In addition, the standard deviation of the voters’ distributions provides a measure of the voter consistency or label ambiguity for a segment (smaller values indicate greater consistency, i.e. less label ambiguity). We explore these annotations in Section  3.2.
@@ 3.1.3 Expert Annotator Qualifications and Training
Unlike the MTurk workers who only viewed 10-second clips for each task, the expert, as a paper author, had previously reviewed the full videos (multiple times) during data annotation and quality control. This exposure provided them with a deeper understanding of the task context and its specific challenges. Additionally, the expert had prior experience coding actions, objects, and errors within the ‘tent’ dataset, further honing their observation and annotation skills relevant to the 10-second clips.
@@
Figure 2: Frequency distributions of struggle labels for each split. The three graphs display the distribution of the four-way struggle labels for the activities Pipes-Struggle (left), Tent-Struggle (middle), and Tower-Struggle (right), respectively.
@@
Figure 4: Comparison of the distributions. The first three graphs on the left display the frequency distribution of the struggle labels including expert annotation (EA), voters mode (Mode), and voters labels distributions (Voters), for the three activities. The three graphs on the right display the histogram with smoothed kernel density estimate (KDE) of the standard deviation (StdDev) of the voters’ labels for the three activities.
@@ 3.1.4 MTurk Annotator Validation
We recognize that struggle determination and its annotation is an ambiguous task, which is part of the challenge and research value for this problem. However, we still want to ensure annotation by crowd-sourcing is as valid as possible and that the votes and MTurkers can be trusted. We use the following assumptions in this ambiguous undertaking: 1) we assume the expert annotator is trustworthy, 2) ambiguity between sides of struggle (e.g. definitely struggling and slightly struggling and similarly for non-struggling) is acceptable and expected, 3) If a Mturker is untrustworthy in one of the assignments, all results by that MTurker is questionable and best to remove. Based on these assumptions we implemented the following validation checks:
@@
  • These 6 validation videos were included as part of the MTurk assignments and for every MTurker we calculate the level of disagreement w.r.t. the validation videos and the expert’s votes. If the MTurker’s votes are on the same side of struggle w.r.t the expert’s vote, i.e. (struggling (3,4) or non-struggling (1,2)) this does not count as a disagreement. The MTurker is rejected if the ratio of disagreement w.r.t. the validation videos is $>0.33$, i.e. three or more out of 6 disagreements for the validation videos.
@@ 3.2 Struggle Statistics
Figure 4 shows both the overall frequencies for each struggle label and the histograms of standard deviations that indicate the label ambiguity for each dataset. The histograms of standard deviations tell us that the Tent-Struggle dataset has the most ambiguous voter labels with a standard deviation between 0.9 and 1. This is followed by the Pipes-Struggle dataset and then the Tower-Struggle dataset, which has the least ambiguous labels. These ambiguities reflect the degree of variation innate to each task with a well-defined task such as Towers of Hanoi having less ambiguity and assembling the non-rigid tent outdoors having the highest. The variation of ambiguity is a further element of interest for struggle-aware systems.
@@ 4 Experiments
Table 3: Results for the classification task. Classification accuracy for binary classification, four-way classification and the four-way to binary classification.
@@
We introduce three tasks for assessing struggle determination. First, struggle classification includes binary and four-way classifications using expert labels. In the four-way classification, we use the original struggle labels ranging from 1 to 4 corresponding to definitely non-struggle to definitely struggle. In binary classification, we binarize the label of struggling versus non-struggling by grouping 1&2 votes and 3&4 votes. Secondly, regression fits to the degree of struggle. Here we treat the struggle level labels from the experts as continuous real numbers ranging from 1 to 4. The third task is label distribution learning. Given the video samples, we train a deep model that can predict the frequency distribution over the four struggle levels while using the frequency distributions of the voters’ annotations for each video segment as the target. In the rest of this section, we present our baseline methods for end-to-end training, followed by the baseline experiment implementation details and results. We train and test our deep models using two Titan X (Pascal) 12GB GPUs or two GTX 1080Ti 11GB GPUs across all the experiments.
4.1 Baseline Methods
We implement three different deep architectures in our experiments: (1) 2D-ConvNet represented by the Temporal Segment Networks (TSN) [58], using the implementation from [63]. (2) 3D-ConvNet represented by the SlowFast Networks [16] based on ResNet50 (SlowFast-R50) [13], and (3) hybrid architectures of 3D-ConvNets and Vision Transformers [8], referred to as VTN-SlowFast-ViT and SlowFast-gMLP. The last two deep architectures implement the SlowFast Networks as the backbone to extract spatial-temporal features from the video sequence and are followed by different attention mechanisms such as a Vision Transformer layer [8] encoder block implemented based on layer [35] (VTN-SlowFast-ViT); or a gMLP layer [30] implemented from [34] (SlowFast-gMLP). Specifically, given a dataset which has a set of $K$ videos $P=\{p_{i},1\leq p_{i}\leq K\}$. We sample $N$ frames from each video sample $p_{i}$ and feed them into the SlowFast backbone $\Phi(\cdot)$, resulting in $N/8$ feature vectors that are then fed into the Vision Transformer encoder or a gMLP layer. This is followed by an MLP head for specific tasks, such as classification, regression, and label distribution learning. We use four-fold cross-validation to train and test these deep models for all tasks. We freeze the parameters in the backbone layers during training to overcome overfitting, which means linear probing for the first two deep architectures and only train the Vision Transformer-based layers for the hybrid architectures.

### v5

@@ 3.1.2 Annotating Struggle
Since the original videos are uniformly trimmed into 10-second segments, we focus on ‘low-level’ struggling cues, which we believe are the most transferable between participants and scenarios. Signs of struggle may include shaky hands, pausing actions, dropping items, or action repetition, however, we note that these are often task dependent. While there is a variety of ways one might ask an observer to estimate struggle, we chose to ask observers to rate video segments using a four-point scale varying from definitely non-struggle to definitely struggle. Table 1 lists the descriptions for each struggle level. The value of a finer, four-point gradation of struggle supports the longer-term goal of assistive systems. For example, the support provided to a person who is ‘slightly’ struggling may consist of a few basic instructions or reminders. However, a person determined to be definitely struggling may require more detailed guidance or even recommendations to stop the task to prevent injury.
This scoring approach allows non-expert observers to indicate the degree of struggle/non-struggle. Note that we forced annotators to choose between struggle and non-struggle instead of allowing a neutral option. This avoided centring strategies on ambiguous judgements and allowed easy binarisation of annotations during early pilot approaches that classed only struggle presence/absence. Throughout our research, we found that indicating the degree of struggle is necessary to address significant challenges in classifying fine-grained features.
As self-reporting levels of struggle can be unreliable and provide only a single judgement, we focus on a balance between expert and naive observers. We collected two distinct types of struggle determination annotations: naive ‘voter labels’ obtained from Amazon Mechanical Turk (MTurk) participants and ‘expert annotations’ (or ‘golden annotation’ (GA)) obtained from the author who collected the video dataset, who repeatedly viewed them during analysis, and annotated action and error labels in the EPIC-Tent dataset. For each 10s video clip, there is only one expert annotation label, and there are multiple voters’ labels that form a frequency distribution. This distribution consists of 15 or 20 voter labels. To extract a single struggle label from this distribution, we take the mode. Since voter labels and expert annotation come from distinct groups of individuals with differing prior knowledge, label distributions can vary across all video segments for each dataset. Consistency between the voters’ modes and expert annotations indicates that the struggle-level annotations of the videos are more reliable. In addition, the standard deviation of the voters’ distributions provides a measure of the voter consistency or label ambiguity for a segment (smaller values indicate greater consistency, i.e. less label ambiguity). We explore these annotations in Section  3.2.
@@ 3.1.3 Expert Annotator Qualifications and Training
Unlike the MTurk workers who only viewed 10-second clips for each task, the expert, as a paper author, had previously reviewed the full videos (multiple times) during data annotation and quality control. For expert annotation, this broader context helped ensure consistency, as the expert worked independently. In contrast, crowd annotators did not require prior task knowledge, but to achieve a relatively consistent struggle annotation, we relied on multiple crowd workers per clip and aggregated their inputs by averaging struggle levels or taking the mode after filtering out outliers. Additionally, the expert had prior experience coding actions, objects, and errors within the ‘tent’ dataset, further honing their observation and annotation skills relevant to the 10-second clips.
@@
Figure 3: Frequency distributions of four-level struggle labels across dataset splits. The three graphs illustrate the distribution of struggle labels (ranging from 1 to 4 on the x-axis) for the activities Pipes-Struggle (left), Tent-Struggle (middle), and Tower-Struggle (right).
@@
Figure 5: Comparison of the distributions. The first three graphs on the left display the frequency distribution of the struggle labels, including expert annotation (EA), voters mode (Mode), and voters labels distributions (Voters), for the three activities. The three graphs on the right display the histogram with smoothed kernel density estimate (KDE) of the standard deviation (StdDev) of the voters’ labels for the three activities.
@@ 3.1.4 MTurk Annotator Validation
We recognize that struggle determination and its annotation is an ambiguous task, which is part of the challenge and research value for this problem. However, we still want to ensure annotation by crowd-sourcing is as valid as possible and that the votes and MTurkers can be trusted. We use the following assumptions in this ambiguous undertaking: 1) we assume the expert annotator is trustworthy, 2) ambiguity between sides of struggle (e.g. definitely struggling and slightly struggling and similarly for non-struggling) is acceptable and expected, 3) If an MTurker is untrustworthy in one of the assignments, all results by that MTurker is questionable and best to remove. Based on these assumptions, we implemented the following validation checks:
@@
  • These 6 validation videos were included as part of the MTurk assignments, and for every MTurker, we calculate the level of disagreement w.r.t. the validation videos and the expert’s votes. If the MTurker’s votes are on the same side of struggle w.r.t the expert’s vote, i.e. (struggling (3,4) or non-struggling (1,2)), this does not count as a disagreement. The MTurker is rejected if the ratio of disagreement w.r.t. the validation videos is >0.33>0.33, i.e. three or more out of 6 disagreements for the validation videos.
@@ 3.2 Struggle Statistics
Figure 5 shows both the overall frequencies for each struggle label and the histograms of standard deviations that indicate the label ambiguity for each dataset. The histograms of standard deviations tell us that the Tent-Struggle dataset has the most ambiguous voter labels with a standard deviation between 0.9 and 1. This is followed by the Pipes-Struggle dataset and then the Tower-Struggle dataset, which has the least ambiguous labels. These ambiguities reflect the degree of variation innate to each task, with a well-defined task, such as Towers of Hanoi, having less ambiguity and assembling the non-rigid tent outdoors having the highest. The variation of ambiguity is a further element of interest for struggle-aware systems. Furthermore, Figure 13 in the Appendix shows the intra-group voter disagreement in the dataset. This figure illustrates additional analysis on the frequency of annotators’ disagreements regarding the binary struggle/non-struggle labels, focusing on the minority group, which has fewer instances of disagreement, ranging from 0 to 0.5. As shown in the figure, video samples with a minority group disagreement frequency of 0.5 constitute only a small proportion of the entire dataset.
@@ 4 Experiments
In this section, we provide baseline results for our struggle determination dataset and structure experiments to answer the following questions: (i) What are the most important modelling approaches for struggle determination? (ii) To what extent can a struggle determination model being trained on one subset generalize to other struggle subsets? (iii) How do different backbone deep architectures impact struggle determination performance? and (iv) What are the most appropriate baseline models for struggle determination?
4.1 Comparison of Struggle Modelling Approaches
Table 3: Comparison of Classification and Regression-to-Classification Accuracy based on MViTv2 [31].
Activity Classification Regression-to-Classification
Binary Accuracy Four-Way Accuracy Binary Accuracy Four-Way Accuracy
Pipes-Struggle 76.68 48.43 75.03 44.71
Tent-Struggle 64.42 37.06 59.86 35.60
Tower-Struggle 84.29 57.42 81.29 33.41
We initially thought about three different modelling approaches to assess struggle determination comprehensively. First, struggle classification includes binary and four-way classifications using expert labels. In the four-way classification, we use the original struggle labels ranging from 1 to 4, corresponding to definitely non-struggle to definitely struggle. In binary classification, we binarize the label of struggling versus non-struggling by grouping 1&2 votes and 3&4 votes. Secondly, regression fits the degree of struggle. Here we treat the struggle level labels from the experts as continuous real numbers ranging from 1 to 4. The third approach is label distribution learning. Given the video samples, we train a deep model that can predict the frequency distribution over the four struggle levels while using the frequency distributions of the voters’ annotations for each video segment as the target. This modelling approach is a supplementary way of making full use of the voters’ annotations, and thus, we will include it in our ablation study (Section 5.3.2).
Compared to struggle classification, struggle regression provides the advantage of modelling struggle levels as continuous real numbers from 1 to 4, capturing subtle variations more effectively. To determine the most suitable modelling approach, we first train a deep model for struggle classification and then train the same model for struggle regression. The regression outputs are then quantized into either binary or four-way struggle classifications, and we evaluate the accuracy rate for comparison. We adopt this regression-to-classification approach because the primary goal in struggle determination is to conclude whether a person is struggling or not, rather than merely predicting a numerical score.
The experimental results are presented in Table 3. We selected MViTv2 [31] for this comparison due to its efficiency in capturing spatiotemporal features as a recent representative transformer-based model. The results show that both binary and four-way struggle classification achieve higher accuracy when the model is trained as a classification task rather than a regression task. This trend holds consistently across all three struggle datasets, suggesting a significant advantage of training the deep model as a classification task. One possible reason is that the regression model predicts a continuous struggle level, and when these scores are quantized into binary or four-way classes, some instances may be misclassified due to threshold selection. We also conducted extensive experiments comparing classification and regression results across different deep models to assess the consistency of this trend (see Appendix Table 9 and 10). The findings consistently indicate that models trained for classification outperform those trained for regression.
4.2 Generalisation Experiments
Table 4: Results show a comparison between training the model on a combined training set versus training on each of the struggle subsets separately. ‘Best Epoch Over Three’ refers to selecting a single epoch that performs relatively best based on the average accuracy across all three validation sets from all three struggle subsets, which is then evaluated individually on each subset. ‘Best Epochs Separately’ refers to choosing the best epochs with the highest accuracy on each subset’s validation set. The results are reported using the mean accuracy across all cross-validation splits.
Dataset Test Sets - Binary Cls. Accuracy (%) Test Sets - Fourway Cls. Accuracy (%)
Training Sets Pipes-Struggle Tent-Struggle Tower-Struggle Pipes-Struggle Tent-Struggle Tower-Struggle
Pipes-Struggle 76.68 49.22 71.56 48.43 21.58 40.79
Tent-Struggle 53.99 64.42 75.54 30.97 37.06 29.04
Tower-Struggle 62.31 47.90 84.29 24.31 25.42 57.42
Combined (Best epoch over three) 77.29 66.71 90.22 48.79 40.26 66.55
Combined (Best epochs Separately) 79.83 69.48 91.74 51.89 41.81 68.81
After identifying struggle classification as a priority modelling approach for struggle determination, we conduct experiments to show the generalisation capability among the three different scenarios: Pipes-Struggle, Tent-Struggle, and Tower-Struggle. We used the same MViTv2 [31] model to conduct the generalisation experiment. We first trained and evaluated the model on the same struggle subset and then conducted zero-shot evaluations on the other two struggle subsets. See Table 4 for the results. Although the struggle classification accuracy dropped when directly evaluated on unseen struggle subsets, some training-testing combinations demonstrate potential generalization. In general, models trained on Pipes-Struggle and Tent-Struggle tend to transfer better to Tower-Struggle, achieving relatively high accuracy. This is followed by models trained on Tent-Struggle and Tower-Struggle transferring to Pipes-Struggle, though the generalization is not as strong. The struggle in the Tent-Struggle subset appears to be the hardest to generalize to, as models trained on Pipes-Struggle and Tower-Struggle achieve the lowest accuracy when tested on Tent-Struggle. This may be due to the struggle patterns in Tent-Struggle being more complex and occurring in an outdoor setting, compared to the other two indoor environments. Additionally, models trained on Tower-Struggle struggle to generalize to the other subsets, especially in four-way classification, where accuracy drops significantly. This may be because the struggle patterns in the Tower-Struggle subset are relatively simple, the task complexity is low, and the number of data samples is the smallest among the three subsets. These findings suggest that certain struggle features are more transferable across subsets, while others remain more dataset-specific.
We also conducted multi-activity learning experiments to train the model using a combination of the training splits of the three struggle subsets and evaluating the model on the validation splits individually. We concatenate the training data to train the MViTv2 [31] model and run three evaluation stages separately to evaluate the accuracy rate on the three subsets. We present results in Table 4, where we report both the binary classification accuracy and the four-way classification accuracy by comparing the combined training results with the baseline results of training separately. The results are reported in two ways: (1) selecting the epochs with the highest accuracy on each validation set individually, which may be different for each set, and (2) selecting a single epoch that provides the best average accuracy across all three validation sets. As shown in Table 4, training the MViTv2 [31] model on combined struggle subsets demonstrates a significant trend of improvement in both binary and four-way classification accuracy, with increases generally ranging from 3% to 11% compared with training on each of the struggle subsets separately. This is not only because the MViTv2 [31] model benefits from having more training data for struggle classification, but also because it may indicate that struggle patterns are potentially transferable across different task-performing activities. We show that the generalisation across datasets is challenging for MViTv2 [31], particularly in a zero-shot setting. Our results indicate that the struggle classification accuracy drops significantly when the model is tested on a different struggle subset. Moreover, MViTv2 [31] requires a large amount of data to generalise well or, more specifically, to surpass the accuracy achieved when trained and tested on individual struggle subsets. This is evident from the bottom two rows of our results, where training on combined subsets leads to better performance on the evaluation test sets.
4.3 Model Comparison on Struggle Classification
4.3.1 Baseline Methods
We compare four different types of deep architectures in our struggle classification experiments: (1) 2D-ConvNet, represented by Temporal Segment Networks (TSN) [68], is implemented using [74]. (2) 3D-ConvNet, represented by SlowFast Networks [16], is based on ResNet50 (SlowFast-R50) [13]. (3) Hybrid architectures, which combine 3D-ConvNets and Vision Transformers [8], include VTN-SlowFast-ViT and SlowFast-gMLP. (4) A Transformer-based model, MViTv2 [31], which has been adopted in the previous experiments. Besides, we included the Video-LLaVA [33] VLM model, evaluated in a zero-shot setting without further fine-tuning, as shown in Appendix C.3.
The two-hybrid architectures, VTN-SlowFast-ViT and SlowFast-gMLP, utilize the SlowFast Networks as the backbone to extract spatial-temporal features from video sequences. These features are then processed by different attention mechanisms: VTN-SlowFast-ViT incorporates a Vision Transformer encoder block [8], implemented using [39], while SlowFast-gMLP applies a gMLP layer [34], implemented using [38].
Specifically, given a dataset containing a set of $K$ videos $P=\{p_{i},1\leq p_{i}\leq K\}$. We sample $N$ frames from each video sample $p_{i}$ and feed them into the SlowFast backbone $\Phi(\cdot)$. This process result in $N/8$ feature vectors, which are then passed through either the Vision Transformer encoder or a gMLP layer. Finally, an MLP head is applied for struggle classification.
We employ four-fold cross-validation to train and evaluate these deep models across all tasks and report the averaged top-1 classification accuracy over all four validation splits. To mitigate overfitting, we freeze the parameters in the backbone layers during training, effectively performing linear probing for TSN [68] and SlowFast-R50 [16]. For the hybrid architectures, only the Vision Transformer-based layers are trained while keeping the backbone frozen.
Table 5: Struggle classification results. Classification accuracy for binary classification, four-way classification and the four-way to binary classification.
@@ 4.3.1 Baseline Methods
MViTv2 [31] 76.68 48.43 72.25 64.42 37.06 58.15 84.29 57.42 82.59
4.3.2 Results and Model Comparison
The struggle classification results are shown in Table 5. The first three rows represent naive baselines: (1) the random performance baseline, which uses the frequency of classes as the probability for random predictions; (2) the majority class baseline, which predicts the majority class and reports its percentage; and (3) the voter’s baseline, where the mode of the voters’ annotations is used to compare with the expert annotations and calculate the averaged accuracy. The experimental results indicate a performance gap between the voter’s baseline accuracy and the best-performing deep models.
Among the evaluated models, TSN [68] and MViTv2 [31] exhibit relatively low accuracy in both binary and four-way classifications. The poor performance of the TSN [68] in struggle classification is probably attributed to the 2D-ConvNet design, which lacks explicit temporal modelling and relies solely on global average pooling (GAP) to aggregate frame-level features. This approach disregards frame order, negatively impacting performance. On the other hand, MViTv2 [31], despite leveraging spatial-temporal dependencies through multi-head self-attention, still underperforms. This may be due to the fact that transformer-based models typically require large amounts of training data to effectively capture useful patterns. This hypothesis is supported by the significant improvement in struggle classification performance when MViTv2 [31] is trained on combined datasets, as shown in Table 4, which also outperforms the struggle classification accuracy of all other deep models shown in Table 5.
In contrast, 3D-ConvNets, represented by the SlowFast [16] and its enhanced variants like VTN-SlowFast-ViT and SlowFast-gMLP, which incorporate more advanced temporal modelling, generally achieve higher accuracy. This suggests that the SlowFast model, as a feature extraction backbone, is more robust for struggle classification, particularly when training data is limited. The struggle classification accuracy is similar across the three models that use SlowFast [16] as the backbone. This suggests that incorporating MLP- or Transformer-based architectures for additional temporal modelling does not significantly improve performance. A likely explanation is that our video-level weak labels for video segments do not require extensive temporal processing beyond the initial spatial-temporal feature extraction. As a result, the potential of the MLP/Transformer layers to enhance accuracy remains limited in this context.
4.4 Implementation Details
We conducted two classification experiments: binary classification, which distinguishes between struggle and non-struggle, and four-class classification, which predicts the level of struggle. In both cases, the model is supervised using the Cross-Entropy (CE) loss. Top-1 accuracy is used as the evaluation metric. During training, we uniformly sample 128 frames from each video and randomly crop the input frames from $256\times 456$ to $256\times 256$. Additionally, a horizontal flip is applied with a probability of 0.5 during training. For the evaluation stage, we perform 5 random crops and ensemble the predictions to improve robustness.
For TSN [68], we implement the BN-Inception base architecture with the average consensus for the output from each segment. We use ImageNet [53] pre-trained features and train only the fully connected layer for classification. To address overfitting, we set the dropout rate to 0.5. The model is trained for 65 epochs using mini-batch stochastic gradient descent with a momentum of 0.9 and a weight decay of $5\times 10^{-4}$. The learning rate starts at 0.0001 and decreases by a factor of $10\times$ at epochs 30 and 60.
For the SlowFast [16] model and its variants, we use the architecture that is based upon 3D-ResNet50 and load parameters pre-trained on Kinetics-400 [28]. During training, we freeze the backbone parameters and train the remaining MLP layers. The same optimizer is used, but the weight decay is set to $1\times 10^{-5}$. We train these models for 70 epochs with a cyclic cosine decay learning rate scheduler [2], where the learning rate starts at 0.001, decreases by a factor of 100 over 50 epochs, and includes two restart intervals of 10 epochs each. During these intervals, the learning rate increases by a factor of 10 at the start and decreases back to the minimum value.
For MViTv2 [31], we implement its small variant to accommodate limited GPU memory. We load parameters pre-trained on Kinetics-400 [28] and fine-tune the model using linear probing to mitigate overfitting. The AdamW optimizer is used with a weight decay of $5\times 10^{-4}$ and a cyclic cosine decay learning rate scheduler. The first 10 epochs serve as a warm-up stage, starting at 0.1 times the base learning rate of 0.001. The learning rate is then gradually decreased by a factor of 0.01 from epoch 10 to epoch 50, followed by two fixed restart intervals where the learning rate decreases from 0.1 times the base learning rate to 0.01 times the base learning rate. The total training duration is 70 epochs. In the generalization experiments, we maintain the same settings but employ a multi-step learning rate scheduler. The base learning rate starts at $1\times 10^{-6}$ and increases by a factor of 10 at epochs 30 and 60, with the models trained for 70 epochs for full parameter fine-tuning.
All experiments are conducted using two Titan X (Pascal) 12GB GPUs or two GTX 1080Ti 11GB GPUs.
5 Ablation Study
Figure 6: Importance of Multiple Frames. The left three bar charts show binary classification results for Pipes-Struggle, Tent-Struggle, and Tower-Struggle, while the right three show four-way classification for the same activities. Each group of bars represents the base model (blue), evaluation with a single frame (orange), and training/evaluation with a single frame (green). Error bars indicate the standard deviation (StdDev) across sampled frames from 25%, 50%, and 75% of the video. See Appendix Table 25 for detailed result numbers. Model abbreviations: TSN [68], SF: SlowFast [16], SFG: SlowFast-gMLP, VSV: VTN-SlowFast-ViT.
Figure 7: Importance of Temporal Ordering. The left three bar charts show binary classification results for Pipes-Struggle, Tent-Struggle, and Tower-Struggle, while the right three show four-way classification for the same activities. Each group of bars represents input frames in the correct order (blue), evaluation with shuffled frames (orange), and training/evaluation with shuffled frames (green). See Appendix Table 26 for detailed result numbers. Model abbreviations: SF: SlowFast [16], SFG: SlowFast-gMLP, VSV: VTN-SlowFast-ViT.
In this section, we investigate the importance of temporal modelling, motion understanding, and alternative training strategies in struggle determination. Specifically, we conduct two series of experiments: (1) validating the use of multiple frames for struggle determination by comparing it with single-frame input and (2) evaluating the importance of temporal ordering in frame sequences for struggle determination. Additionally, we explore alternative training strategies by presenting baseline results for struggle regression and struggle label distribution learning. We also compare different deep learning models to identify the most effective architecture for struggle determination under these two training approaches.
5.1 Multiple Frames
To demonstrate the importance of using multiple frames, we conducted experiments comparing a single-frame input to the original multi-frame setting.
First, we initialized the model with parameters trained on 128 frames, which is the standard setting for our experiments. We then tested the model using only one frame from each video sequence. The selected frame was taken from three different positions within each video segment—at 25%, 50%, and 75% of its duration. Since 3D-ConvNets require a minimum number of input frames, we duplicated the selected frame to meet these requirements: 8 frames for SlowFast [16] and 128 frames for SlowFast-gMLP and VTN-SlowFast-ViT.
To ensure consistency between training and testing, we conducted an additional experiment where we trained the model using randomly selected single frames from each video. During testing, we applied the same single-frame selection strategy as before, choosing frames from 25%, 50%, and 75% of each video segment.
The results of these two approaches are shown in Figure 6. We report the average test accuracy for single-frame inputs across the three proportions. Across all tested models and activities, we observed a significant drop in accuracy when using only one frame. Since our dataset consists of randomly trimmed 10-second video segments, accuracy values for frames sampled from 25%, 50%, and 75% are expected to be similar, leading to a small standard deviation (StdDev) in accuracy, as reflected in the bar plots.
Even when the models were trained using a single frame, the test accuracy remained notably lower compared to the multi-frame setting, though slightly higher than testing on a single frame without training adjustments. This suggests that a single frame does not provide sufficient visual information for deep learning models to reliably determine whether a person is struggling.
Therefore, effectively capturing the useful temporal motion features is crucial for our struggle determination task. A sequence of frames can typically reveal anomalous hand motions such as ‘unsmooth’ motions, stopping, and repeating that indicate the person did not have things under control. A single frame cannot represent these dynamic features.
5.2 Temporal Ordering